OpenStudy (kcnkinoie):

Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form.

OpenStudy (kcnkinoie):

x – 4 = –1(y – 2) x – 2 = –1(y – 4) y – 2 = –1(x – 4) y – 4 = –1(x – 2)

OpenStudy (zyi6):

Since we know what the ordered pair is, and we have the slope, we can solve this very quickly: Point slope form: y - y1 = m(x - x1) So, we just substitute our values in: y - 4 = -1(x - 2) is the answer!